### v1

@@ 3 Coordination Reduces RoS Violation
[…]
Letting ${\varepsilon}=\frac{\Delta T}{2N}$ proves the lemma. ∎
Lemmas 1 and 2 together prove Theorem 3.1.
4 Coordination Increases Bidders’ Value
In this section, we show that coordination can increase the total value of auto-bidders, compared to independent bidding.
We will focus on a specific auto-bidding algorithm: the mirror-descent algorithm in feng2023online. It is the state-of-the-art algorithm for value maximization under RoS constraint, with near-optimal performances for the bidders, achieving $O(\sqrt{T})$ regrets. The algorithm maintains a parameter $\lambda_{i,t}$ for each bidder (which corresponds to the Lagrange multiplier of the dual optimization problem), so the bidder overbids by $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$ at each round. After observing a utility feedback, the bidder adjusts the parameter $\lambda_{i,t}$ in the opposite direction of the utility (i.e., positive utility decreases $\lambda_{i,t}$). The full algorithm is given in Algorithm 3.
Initialize multiplier $\lambda_{i,1}=1$, learning rate $\alpha=1/\sqrt{T}$;
for $t=1,2,\dots,T$ do
   Observe value $v_{i,t}$ and set bid $b_{i,t}=(1+\frac{1}{\lambda_{i,t}})v_{i,t}$;
   Observe allocation $x_{i,t}$ and payment $p_{i,t}$, compute utility $g_{i,t}=v_{i,t}\cdot x_{i,t}-p_{i,t}$;
   Update the multiplier: $\lambda_{i,t+1}=\lambda_{i,t}\exp(-\alpha g_{i,t})$;
end for
ALGORITHM 3 Auto-bidding algorithm in feng2023online
For Algorithm 3, we show that coordinated auto-bidders obtain higher total value than independent auto-bidders. Let
$V_{t}^{C}=\sum_{i=1}^{N}v_{i,t}x^{C}_{i,t},\quad V_{t}^{I}=\sum_{i=1}^{N}v_{i,t}x^{I}_{i,t}$
be the total values obtained by the $N$ bidders at round $t$ in the coordinated case and independent case, respectively.
Theorem 4.1(Coordination Increases Total Value)
Assume 1, and suppose bidders run Algorithm 3. As $T\to\infty$, in expectation, the bidders’ total value under coordinated bidding is larger than that under independent bidding:
$\lim_{T\to\infty}\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{C}_{t}-\frac{1}{T}\sum_{t=1}^{T}V^{I}_{t}\Big]\geq 0.$
To prove Theorem 4.1, we first reduce the $N$-bidder coordinated scenario to a single-bidder problem where the single bidder’s value is the maximum among the coalition members. Using this reduction, we prove an interesting fact: under Assumption 1, the multipliers $\lambda_{i,t}^{C}$ of coordinated bidders converge to 0 in the limit as $t\to\infty$, meaning that the coordinated bidders will eventually submit arbitrarily high bids. This implies that the coordinated coalition will obtain a higher value than the independent case. We present the full proof below.
4.1 Proof of Theorem 4.1
@@ 4 Coordination Increases Bidders’ Value
Since the $N$-bidder coordinated scenario is equivalent to a single-bidder scenario, we can now analyze the single-bidder scenario. We present two useful lemmas regarding the expected utility function $G(\lambda)$ of the single bidder.
Lemma 3
$G(\lambda)$ is monotonically increasing in $\lambda$.
Proof
See Appendix 0.A. ∎
Lemma 4
Under Assumption 1, for any $\lambda>0$, $G(\lambda)\geq\Delta>0$.
Proof
By Lemma 3, $G(\lambda)$ is increasing in $\lambda$. So,
@@
$\geq\mathbb{E}\big[(v_{(N-1)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]\geq\Delta>0$
@@
Intuitively, Lemma 4 shows that the expected utility obtained by the highest-value coalition bidder is always positive (by a margin of $\Delta>0$). This means that the bidder’s multiplier $\lambda_{t}$, following mirror descent updates, will decrease in expectation every round and eventually converge to 0. We formalize this intuition below.
@@
Lemma 5
Under Assumption 1, for any $i\in[N],t\in[T]$, with probability at least $1-\exp\big(-\frac{\Delta^{2}t}{32B^{2}N^{2}}\big)$, $\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\big(-\frac{\Delta t}{2N\sqrt{T}}\big)$.
Proof
According to Algorithm 3, $\lambda_{i,t+1}^{C}$ is equal to
$\lambda_{i,t+1}^{C}=\lambda_{i,1}\exp\Big(-\alpha\sum_{t^{\prime}=1}^{t}g_{i,t^{\prime}}\Big)$
where each $g_{i,t^{\prime}}$ is an unbiased estimator of bidder $i$’s expected utility $G_{i}^{C}(\lambda_{i,t^{\prime}})=\mathbb{E}[g_{i,t^{\prime}}]$ at round $t^{\prime}$. If bidder $i$ is not the highest-value bidder in round $t^{\prime}$, then $g_{i,t^{\prime}}=0$, meaning that the bidder’s multiplier $\lambda_{i,t^{\prime}}^{C}$ is not updated in that round. According to Observation 1 and Lemma 4, bidder $i$’s expected utility is at least
@@
$\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\Big(-\alpha\big(t\cdot\frac{\Delta}{N}-{\varepsilon}\big)\Big)=\lambda_{i,1}\exp\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big).$
@@
$V_{(N)}(\lambda)=\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Big[v_{(N)}\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\big]\Big].$
@@
$V_{(N)}(0^{+})=\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$
exists. Note that the expected total value of $N$ coordinated bidders with multipliers $\lambda_{1}^{C},\ldots,\lambda_{N}^{C}$ is
$V^{C}(\lambda_{1}^{C},\ldots,\lambda_{N}^{C})=\mathbb{E}\Big[v_{i^{*}}\!\cdot\mathbb{I}\big[(1+\tfrac{1}{\lambda_{i^{*}}})v_{i^{*}}>d^{O}\big]\Big]=\mathbb{E}\big[V_{(N)}(\lambda_{i^{*}}^{C})\big]$
where $i^{*}=\operatorname*{arg\,max}_{i\in[N]}v_{i}$. This means that the expected average total value of the coordinated bidders is
$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{C}\Big]=\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i^{*}_{t},t}^{C})\Big]$ (3)
where $i^{*}_{t}\in\{1,\dots,N\}$ is the highest-value bidder in round $t$.
Let $V^{I}(\lambda_{1},\ldots,\lambda_{N})$ be the expected total value of $N$ independent bidders with multipliers $\lambda_{1},\ldots,\lambda_{N}>0$:
$V^{I}(\lambda_{1},\ldots,\lambda_{N})=\mathbb{E}_{\bm{v}\sim F,d^{O}\sim D}\Big[\sum_{i=1}^{N}v_{i}x_{i}(\bm{b}^{I})\Big]$
where $\bm{b}^{I}=\big((1+\tfrac{1}{\lambda_{1}})v_{1},\ldots,(1+\tfrac{1}{\lambda_{N}})v_{N},d^{O}\big)$.
Lemma 6
$V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$, for any $\lambda_{1},\ldots,\lambda_{N}>0$.
Proof
Pick $\lambda_{\varepsilon}<\min\{\lambda_{1},\dots,\lambda_{N}\}$. If some bidder $i\in[N]$ wins the item under independent bidding, then their bid $b_{i}^{I}=(1+\frac{1}{\lambda_{i}})v_{i}$ is larger than the outside bid $d^{O}$. In that case, suppose the highest-value bidder uses multiplier $\lambda_{\varepsilon}$, then their bid is $(1+\frac{1}{\lambda_{\varepsilon}})v_{(N)}\geq(1+\frac{1}{\lambda_{i}})v_{i}=b_{i}^{I}>d^{O}$, so the highest-value bidder also wins in the single-bidder scenario. Note that the value obtained by the highest-value bidder is at least $v_{i}$, so the expected value $V_{(N)}(\lambda_{\varepsilon})\geq V^{I}(\lambda_{1},\ldots,\lambda_{N})$. Letting $\lambda_{\varepsilon}\to 0$ proves the lemma. ∎
Proof(Proof of Theorem 4.1)
By Lemma 5, with constant $c\coloneqq\frac{\Delta^{2}}{32B^{2}N^{2}}$, for any fixed $i$ and $t\geq 1$, the coordinated multiplier $\lambda_{i,t+1}^{C}$ satisfies
$\Pr\!\left[\lambda_{i,t+1}^{C}\leq\lambda_{i,1}\exp\!\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big)\right]\;\geq\;1-e^{-ct}.$ (4)
@@
$m_{T}\;\coloneqq\;\left\lceil\frac{2N\sqrt{T}}{\Delta}\,\log\!\Big(\frac{\lambda_{i,1}}{\delta}\Big)\right\rceil~=~O\Big(\sqrt{T}\log\frac{1}{\delta}\Big),$
@@
$\lambda_{i,1}\exp\!\Big(-\frac{t\Delta}{2N\sqrt{T}}\Big)\leq\lambda_{i,1}\exp\!\Big(-\frac{m_{T}\Delta}{2N\sqrt{T}}\Big)\leq\delta.$
@@
On this event, for all $t\geq m_{T}$ we have $\lambda^{C}_{i^{*}_{t},t}\leq\delta$, hence by Eq. (5),
@@
$\mathbb{E}\bigg[\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\bigg]\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon+B\,p_{T}$
$\;\leq\;\frac{O(\sqrt{T}\log\tfrac{1}{\delta})}{T}\,B+\varepsilon+BNTe^{-cO(\sqrt{T}\log\tfrac{1}{\delta})}~\leq~3{\varepsilon}$
@@
Combining with Lemma 6, we conclude that
$V_{(N)}(0^{+})\;\geq\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V^{I}(\lambda_{1,t}^{I},\ldots,\lambda_{N,t}^{I})\Big]\;=\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{I}\Big],$

### v2

@@ B.2 Boundary case $\lambda_{\star}=0$
Proof.
In coordination, only $i_{t}^{\star}$ bids against $d_{t}^{O}$, so the realized value in round $t$ is
$V_{T}^{C,A}=v_{(N),t}\mathbb{I}[\bigl(1+1/\lambda_{i_{t}^{\star},t}\bigr)v_{(N),t}>d_{t}^{O}].$
Condition on $H_{t-1}$. The fresh randomness $(v_{(N),t},d_{t}^{O})$ is independent of $H_{t-1}$, hence
$\mathbb{E}[V_{T}^{C,A}\mid H_{t-1}]=V_{(N)}(\lambda_{i_{t}^{\star},t}).$
Therefore
$\mathbb{E}[V_{T}^{C,A}]=\mathbb{E}\Bigl[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i_{t}^{\star},t})\Bigr].$
Fix $\varepsilon>0$. By definition of the right-limit $V_{(N)}(0^{+})$ and boundedness $0\leq V_{(N)}(\cdot)\leq B$, there exists $\delta\in(0,1)$ such that
$|V_{(N)}(\lambda)-V_{(N)}(0^{+})|\leq\varepsilon\qquad\forall\lambda\in(0,\delta].$
Hence
$\Bigl|\mathbb{E}[V_{T}^{C,A}]-V_{(N)}(0^{+})\Bigr|\leq\varepsilon+2B\cdot\mathbb{E}\Bigl[\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}[\lambda_{i_{t}^{\star},t}>\delta]\Bigr].$
The second term vanishes by Lemma B.3, so $\limsup_{T}|\mathbb{E}[V_{T}^{C,A}]-V_{(N)}(0^{+})|\leq\varepsilon$. Since $\varepsilon$ is arbitrary, the limit is 0. ∎
Lemma B.5 (Independent value is always $\leq V_{(N)}(0^{+})$).
For any fixed multipliers $(\lambda_{1},\dots,\lambda_{N})\in(0,\infty)^{N}$, let $V^{I}(\lambda_{1},\dots,\lambda_{N})$ denote the independent-bidding expected total value in one round:
$V^{I}(\lambda_{1},\dots,\lambda_{N}):=\mathbb{E}\Bigl[\sum_{i=1}^{N}v_{i}\,\mathbb{I}[i\text{ wins under bids }b_{i}=(1+1/\lambda_{i})v_{i}]\Bigr].$
for all $(\lambda_{1},\dots,\lambda_{N})$,
$V^{I}(\lambda_{1},\dots,\lambda_{N})\leq V_{(N)}(0^{+}).$
Consequently, for the dynamic independent process,
$\mathbb{E}[V_{T}^{I,A}]\leq V_{(N)}(0^{+})\qquad\forall T,\quad\text{and hence}\quad\limsup_{T\to\infty}\mathbb{E}[V_{T}^{I,A}]\leq V_{(N)}(0^{+}).$
Proof.
Fix $(\lambda_{1},\dots,\lambda_{N})$ and pick any $\lambda_{\varepsilon}<\min_{i}\lambda_{i}$. Consider the single-bidder scenario where a bidder with value $v_{(N)}:=\max_{i}v_{i}$ bids with multiplier $\lambda_{\varepsilon}$ against the outside bid $d^{O}$ (no internal competition). Its bid is $b^{\text{single}}=(1+1/\lambda_{\varepsilon})v_{(N)}$.
For any realization $(v_{1},\dots,v_{N},d^{O})$, if the coalition loses under independent bidding, then the independent value is 0 and the single-bidder value is $\geq 0$, so the inequality holds. If some bidder $i$ wins under independent bidding, then its bid satisfies $b_{i}=(1+1/\lambda_{i})v_{i}>d^{O}$ and $b_{i}\geq\max_{j\neq i}b_{j}$. Since $\lambda_{\varepsilon}<\lambda_{i}$, we have $(1+1/\lambda_{\varepsilon})\geq(1+1/\lambda_{i})$, hence
$b^{\text{single}}=(1+1/\lambda_{\varepsilon})v_{(N)}\geq(1+1/\lambda_{i})v_{i}=b_{i}>d^{O}.$
Thus the single bidder wins whenever independent wins. Moreover, its realized value equals $v_{(N)}\geq v_{i}$, so the realized single-bidder value is pointwise at least the realized independent coalition value. Taking expectations yields
$V_{(N)}(\lambda_{\varepsilon})\ \geq\ V^{I}(\lambda_{1},\dots,\lambda_{N}).$
Letting $\lambda_{\varepsilon}\downarrow 0$ and using monotone convergence/bounded convergence gives $V_{(N)}(0^{+})\geq V^{I}(\lambda_{1},\dots,\lambda_{N})$.
For the dynamic process, condition on $H_{t-1}$ to fix $(\lambda_{1,t},\dots,\lambda_{N,t})$; then $\mathbb{E}[V_{T}^{I,A}\mid H_{t-1}]=V^{I}(\lambda_{1,t},\dots,\lambda_{N,t})\leq V_{(N)}(0^{+})$. Average over $t$ and take expectations. ∎
Theorem B.1.
$\liminf_{T\to\infty}\ \mathbb{E}[V_{T}^{C,A}-V_{T}^{I,A}]\ \geq\ 0.$
Proof.
By Lemma B.4, $\lim_{T\to\infty}\mathbb{E}[V_{T}^{C,A}]=V_{(N)}(0^{+})$. By Lemma B.5, $\limsup_{T\to\infty}\mathbb{E}[V_{T}^{I,A}]\leq V_{(N)}(0^{+})$. Therefore,
$\liminf_{T\to\infty}\mathbb{E}[V_{T}^{C,A}-V_{T}^{I,A}]\geq\lim_{T\to\infty}\mathbb{E}[V_{T}^{C,A}]-\limsup_{T\to\infty}\mathbb{E}[V_{T}^{I,A}]\geq V_{(N)}(0^{+})-V_{(N)}(0^{+})=0.$
B.3 Proof of Theorem 4.2
@@
Since the $N$-bidder coordinated scenario is equivalent to a single-bidder scenario, we can now analyze the single-bidder scenario.
In addition to Lemma B.1, we present the following useful lemma regarding the expected utility function $G(\lambda)$ of the single bidder, under Assumption 1.
Lemma B.6.
Under Assumption 1, $G(\lambda)\geq\Delta>0$.
Proof.
By Lemma B.1, $G(\lambda)$ is increasing in $\lambda$. So,
@@
$\geq\mathbb{E}\big[(v_{(N-1)}-d^{O})_{+}\big]-\mathbb{E}\big[(d^{O}-v_{(N)})_{+}\big]$
$\geq\Delta>0$
@@
Intuitively, Lemma B.6 shows that the expected utility obtained by the highest-value coalition bidder is always positive (by a margin of $\Delta>0$). This means that the bidder’s multiplier $\lambda_{t}$, following mirror descent updates, will decrease in expectation every round and eventually converge to 0. We formalize this intuition below.
@@
Lemma B.7.
Under Assumption 1, for any $i\in[N],t\in[T]$, with probability at least $1-\exp\big(-\frac{\Delta^{2}t}{32B^{2}N^{2}}\big)$, $h^{\prime}(\lambda_{i,t+1}^{C})\leq h^{\prime}(\lambda_{i,1})-\frac{\Delta t}{2N\sqrt{T}}$.
Proof.
For simplicity, let $y_{i,t}:=h^{\prime}(\lambda_{i,t})$. According to Algorithm 3 and equation (7),
$y_{i,t+1}=y_{i,t}-\alpha g_{i,t}$ (25)
where each $g_{i,t^{\prime}}$ is an unbiased estimator of bidder $i$’s expected utility $G_{i}^{C}(\lambda_{i,t^{\prime}})=\mathbb{E}[g_{i,t^{\prime}}]$ at round $t^{\prime}$. If bidder $i$ is not the highest-value bidder in round $t^{\prime}$, then $g_{i,t^{\prime}}=0$, meaning that the bidder’s multiplier $\lambda_{i,t^{\prime}}^{C}$ is not updated in that round. According to Observation 2 and Lemma B.6, bidder $i$’s expected utility is at least
@@
$y_{i,t+1}^{C}\leq y_{i,1}-\alpha\left(t\cdot\frac{\Delta}{N}-{\varepsilon}\right)=y_{i,1}-\frac{t\Delta}{2N\sqrt{T}}.$
@@
$V_{(N)}(\lambda)~:=~\mathbb{E}_{v_{(N)}\sim F_{(N)},d^{O}\sim D}\Bigl[v_{(N)}\mathbb{I}\left[(1+\tfrac{1}{\lambda})v_{(N)}>d^{O}\right]\Bigr]$ (26)
@@
$V_{(N)}(0^{+})~:=~\lim_{\lambda\to 0^{+}}V_{(N)}(\lambda)$
exists. Note that the expected total value of the $N$ coordinated bidders with multipliers $\lambda_{1}^{C},\ldots,\lambda_{N}^{C}$ is
$V^{C}(\lambda_{1}^{C},\ldots,\lambda_{N}^{C})~=~\mathbb{E}\Big[v_{i^{*}}\mathbb{I}\big[(1+\tfrac{1}{\lambda_{i^{*}}})v_{i^{*}}>d^{O}\big]\Big]~=~\mathbb{E}\big[V_{(N)}(\lambda_{i^{*}}^{C})\big]$
where $i^{*}=\operatorname*{arg\,max}_{i\in[N]}v_{i}$. This means that the expected average total value of the coordinated bidders during $T$ rounds is
$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{C,A}\Big]~=~\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda_{i^{*}_{t},t}^{C})\Big]$ (27)
where $i^{*}_{t}\in[N]$ is the highest-value bidder in round $t$.
Lemma B.8.
For any coordination mechanism $\mathcal{G}$ and any round $t$,
$\mathbb{E}\!\left[V_{t}^{\mathcal{G}}\right]~\leq~V_{(N)}(0^{+})~=~\mathbb{E}_{v_{(N)}\sim F_{(N)}}\big[v_{(N)}\big].$
Consequently, for every $T\geq 1$,
$\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{\mathcal{G}}\Big]~\leq~V_{(N)}(0^{+}).$
Proof.
Fix a round $t$ and write $v_{(N),t}\coloneqq\max_{i\in[N]}v_{i,t}$. In a single-item auction, the coalition’s realized value is either 0 (if the outside bid wins) or equals the realized value of the winning coalition bidder, which is always at most $v_{(N),t}$. Hence $V_{t}^{\mathcal{G}}\leq v_{(N),t}$ almost surely, and taking expectations gives $\mathbb{E}[V_{t}^{\mathcal{G}}]\leq\mathbb{E}[v_{(N),t}]$. Finally, since bids and $d^{O}$ are bounded, $(1+\tfrac{1}{\lambda})v_{(N)}\to+\infty$ as $\lambda\downarrow 0$, so the single bidder wins with probability $1$ in the limit and thus $V_{(N)}(0^{+})=\mathbb{E}[v_{(N)}]$. ∎
Proof of Theorem 4.2.
By Lemma B.7, with constant $c\coloneqq\frac{\Delta^{2}}{32B^{2}N^{2}}$, for any fixed $i$ and $t\geq 1$, the coordinated multiplier $\lambda_{i,t+1}^{C}$ satisfies
$\Pr\!\left[h^{\prime}(\lambda_{i,t+1}^{C})\leq h^{\prime}(\lambda_{i,1})-\frac{t\Delta}{2N\sqrt{T}}\right]\;\geq\;1-e^{-ct}.$ (28)
@@
Since $h^{\prime}$ is strictly increasing and $\lim_{\lambda\downarrow 0}h^{\prime}(\lambda)=$-$\infty$, we can define $b(\delta):=(h^{\prime})^{-1}(\delta)$.
@@
$m_{T}\;\coloneqq\;\left\lceil\frac{2N\sqrt{T}}{\Delta}(h^{\prime}(\lambda_{i,1})-b(\delta))\right\rceil~=~O\Big(\sqrt{T}(-b(\delta))\Big),$
@@
$h^{\prime}(\lambda_{i,1})-\frac{t\Delta}{2N\sqrt{T}}\leq h^{\prime}(\lambda_{i,1})-\frac{m_{T}\Delta}{2N\sqrt{T}}\leq b(\delta).$
@@
On this event, for all $t\geq m_{T}$ we have $h^{\prime}(\lambda_{i_{t}^{*},t}^{C})\leq b(\delta)$, which is equivalent to $\lambda^{C}_{i^{*}_{t},t}\leq\delta$, hence by Eq. (29),
@@
$\mathbb{E}\bigg[\Big|V_{(N)}(0^{+})-\frac{1}{T}\sum_{t=1}^{T}V_{(N)}(\lambda^{C}_{i^{*}_{t},t})\Big|\bigg]$
$\;\leq\;\frac{m_{T}}{T}\,B+\frac{T-m_{T}}{T}\,\varepsilon+B\,p_{T}$
$\;\leq\;\frac{O(\sqrt{T}(-b(\delta)))}{T}\,B+\varepsilon+BNTe^{-cO(\sqrt{T}(-b(\delta)))}~\leq~3{\varepsilon}$
@@
Combining with Lemma B.8, we conclude that
$V_{(N)}(0^{+})\;\geq\;\mathbb{E}\Big[\frac{1}{T}\sum_{t=1}^{T}V_{t}^{\mathcal{G}}\Big],$